In a region, 31% of people are smokers, 19% of smokers develop lung cancer and 2% of non-smokers develop lung cancer. What is the probability that a person chosen at random willdevelop lung cancer; i.e. what percentage of this region will suffer from lung cancer?
Answer #1

31% of people are smokers, 19% of smokers develop lung cancer and 2% of non-smokers develop lung cancer.

So the probability that one person chosen at random will develop lung cancer = (0.31)(0.19)+(0.69)(0.02) = 0.0589+0.0138 = 0.0727

Hence the required probability is 0.0727 i.e 7.27% of the region will suffer from lung cancer.